. If you earned an average of
$25,000 over your working life and you retire after 2005 at
age 62, 63, or 64, then your annual Social Security benefit
will be $7000, $7500, or $8000, respectively.
There is a linear equation that gives the annual benefit b in terms of age a for
these three years. Find the equation.

Answer by stanbon(75887)

If you earned an average of
$25,000 over your working life and you retire after 2005 at
age 62, 63, or 64, then your annual Social Security benefit
will be $7000, $7500, or $8000, respectively.
There is a linear equation that gives the annual benefit b in terms of age a for
these three years. Find the equation.
-----
Pick two points:
(0,7000),(1,7500)
---
slope: 500/1
---
Form: y = mx + b
7500 = 500*1+b
b = 7000
----------------------
Equation:
Benefit = 500(x)+7000 where "x" is the number of years after age 62
--------------------
Cheers,
Stan H.